This week, Super-Save Supermarket lowered the price of apples from $1 to 90 cents per pound. The quantity of apples sold last week was 200 pounds. This week, the quantity sold was 250 pounds. Calculate the price elasticity of demand. Is it elastic, inelastic, or unitary elastic? What happens to total revenue? 


1
Expert's answer
2021-01-27T08:54:29-0500

"PED=\frac{change in Q.D} {change in price}"


Calculating percentage


The price decreases from $1 to $0.90. Therefore % change = "\frac{-0.1} {1} = -0.1 (-10)"


-0.1 = -10% (-0.1 "\u00d7" 100)


Quantity increased by 50/200 = 0.25 (25%)


Therefore PED = "\frac{25} {-10}"


Therefore PED = 2.5


PED is eelelastic

The revenue will increase